If x,y are integral solutions of 2x- 3xy - 2y2 = 7, then value of  mod(x + y) is

Grade:12

4 Answers

manjeet
10 years ago
the person who have asked this is a dumb
Rahul Mishra
9 years ago
The equation can be written as 2x^2-4xy+xy-2y^2=7 which can be written as (x-2y)(2x+y)=7
,now 7 is prime number means it has only two factors 7 or 1 ,so either x-2y=1 and 2x+y=7 or 2x+y=1 and x-2y=7,by solving these we get x-2y=1 and 2x+y=7 is correct because it gives both x and y as integer and answer is |x+y|=4..................thanks to me
kartheeek
9 years ago
MOD(X+Y)=4 IS One solution and the other is  x-2y=-1 and 2x+y=-7 and its permutation then we get x=-3 and y=-1 then mod(x+y)=mod(-3+-1)=mod(-4)=4
Khushi sharma
6 years ago
|x+y|=4 simply you`ll get the answer ................................?.......... You will get x = 3 or -3 and value of y =1 or -1 but while taking the whole modulus we will get a positive value which is equal to four .....